# Genetic Instruct: Scaling up Synthetic Generation of Coding Instructions for Large Language Models

### v1

@@ 4.1 Experimental Settings
Table 1: Comparison of models fine-tuned with Genetic-Instruct with other baselines. Cells marked with $\times$ indicate result could not be calculated since the model is not available. The column labelled as OpenAI indicates whether the synthetic data is generated with any of the OpenAI’s models like ChatGPT. Best results for each base model are bolded. Cells marked with ✝ are evaluated on the original 500 un-sanitized MBPP dataset and aren’t directly comparable to the MBPP’s result calculated by EvalPlus.
Size Base Model Generation Algorithm OpenAI HumanEval HumanEval+ MBPP MBPP+
47B Mixtral-8x7B - N 40.2 34.3 60.7 50.1
- GPT-4-Turbo - Y 85.4 81.7 83.5 70.7
7B DeepSeek-Base WaveCoder Y 64.0 $\times$ 62.8 ✝ $\times$
7B CodeLlama-Python WaveCoder Y 48.1 $\times$ 47.2 $\times$
7B CodeLlama-Python WizardCoder Y 55.5 45.9 59.9 49.1
34B CodeLlama-Python WizardCoder Y 73.2 64.6 73.2 59.9
7B DeepSeek-Base - N 47.4 39.5 60.6 52.1
Genetic-Instruct N 57.9 51.8 70.4 59.1
[…]
CodeLlama-Instruct N 34.8 30.2 * 44.4 38.7 *
Genetic-Instruct N 50.6 47.6 65.2 55.9
34B CodeLlama-Python - N 53.7 46.2 * 56.2 49.3 *
CodeLlama-Instruct N 41.5 33.7 * 57.0 50.5 *
Genetic-Instruct N 59.1 53.7 67.7 57.1
70B CodeLlama-Python - N 57.3 51.3 * 65.6 53.4 *
CodeLlama-Instruct N 67.8 58.3 * 62.2 51.8 *
Genetic-Instruct N 57.9 55.5 69.2 58.6
We trained the base models using the AdamW optimizer (Kingma and Ba, 2015) with a learning rate of 5e-6 and zero weight decay for two epochs. The learning rate scheduler employed is CosineAnnealing with 100 warmup steps. The final model is obtained by averaging the last four checkpoints, which are saved every 500 steps. All models are trained using tensor parallelism, and BF16 precision is utilized to accelerate the training process.
@@
For large throughput inference with large effective batch sizes, we employ vLLM (Kwon et al., 2023) as our inference engine. Nucleus sampling (Holtzman et al., 2020) is used for decoding, with a temperature of 0.95 for the Instructor-LLM and a temperature of 1.0 for the Coder-LLM and Judge-LLM. To enhance GPU utilization and expedite the generation process, we implement both inter-node and intra-node parallelizations. Specifically, we use 20 colonies for each generation in inter-node parallelization and 10 sub-colonies per colony in intra-node parallelization. A maximum sequence length of 1024 is set for all LLMs to optimize the generation speed and memory allocation.
@@
For the mutation operation, a batch size of 100 per sub-colony is employed. In the crossover operation, a batch size of 10 per sub-colony is used, as we find that the model approximately generates 10 unique instructions on average per sample, resulting in a similar number of generated samples per batch. We randomly choose to perform either crossover or mutation on each batch. The Instructor-LLM uses 3-shot samples for in-context learning and is instructed to generate up to 20 new instructions.
The deduplication process is conducted using the MinHash algorithm from the TextDedup library [^2] https://github.com/ChenghaoMou/text-dedup with a 5-gram overlap, 250 permutations, and a threshold of 0.7. The synthetic dataset generated by our proposed algorithm, Genetic-Instruct, comprises approximately 1.3 million coding instructions, which are divided into different subsets for evaluation. More details on the generated dataset can be found in Appendices F and G.
Table 2: Comparing the effectiveness of different operations in the Genetic-Instruct algorithm. Model CodeLlama-Python 7B has been used as the base model in these experiments. Bolded cells indicate best score for a given dataset size.
Population Size Generation Algorithm HumanEval HumanEval+ MBPP MBPP+
- CodeLlama-Python 7B 38.4 35.2 * 47.6 40.5 *
Cross-Over Only 47.0 41.5 55.4 44.4
2K Mutation Only 40.2 34.8 60.4 47.1
Genetic-Instruct 46.3 41.5 56.4 45.4
Cross-Over Only 48.2 45.7 59.1 51.6
75K Mutation Only 43.9 39.0 63.2 51.9
Genetic-Instruct 49.4 45.7 62.7 52.9
Cross-Over Only 51.7 47.6 62.2 52.9
250K Mutation Only 47.6 42.1 62.7 52.1
Genetic-Instruct 51.8 47.6 63.9 54.6
Table 3: Ablation study on the effect of the generator model on data generation quality. We did the evaluations on two base models: DeepSeek Coder 6.7B and 33B. The first three rows are the generation models used for data synthesis.
Model Dataset Size HumanEval HumanEval+ MBPP MBPP+
Mixtral 8x7B - 40.2 34.3 60.7 50.1
Mixtral 8x22B - 75.3 71.1 71.9 63.6
Llama 3 70B - 77.1 71.2 80.3 67.8
DeepSeek Coder 6.7B Base 47.4 39.5 60.6 52.1
Mixtral 8x7B 900K 54.9 51.8 71.4 60.4
Mixtral 8x22B 377K 66.5 62.8 72.2 60.4
Mixtral 8x22B 873K 68.3 66.5 69.9 59.9
Mixtral 8x22B 1.3M 70.7 67.7 75.4 63.7
Llama 3 8B 400K 63.4 57.9 70.2 58.6
Llama 3 8B 1.2M 64.6 61.0 69.9 57.1
Llama 3 70B 767K 66.5 62.8 72.2 61.4
Llama 3 70B 1.8M 70.7 66.5 75.4 63.2
DeepSeek Coder 33B Base 51.2 44.5 65.4 56.8
Llama 3 70B 767K 70.1 65.9 77.9 64.7
Llama 3 70B 1.8M 73.8 68.9 77.7 64.7
@@ 4.2 Performance Evaluation
In this section, we evaluate the effectiveness of our proposed approach to generate synthetic SFT samples that can enhance the coding capabilities of LLMs. We used CodeLLama-Python-7B (Rozière et al., 2024) and DeepSeek-Coder-6.7B (Guo et al., 2024) as the base models and employed Mixtral 8X7B (Jiang et al., 2024) as the Instructor-LLM, Coder-LLM, and Judge-LLM. The results, along with the performance of other coding models, are reported in Table 1. For cases where no generation algorithm is used, the results of the base models are reported. All results are obtained using greedy decoding with the pass@1 metric. The base models are fine-tuned with a subset of 405K samples from all Genetic-Instruct’s generations, while the WizardCoder and WaveCoder models are trained with datasets of 78K and 20K samples, respectively.
The results show that training models with Genetic-Instruct significantly improves performance on all benchmarks compared to the base models. This demonstrates the effectiveness of the Genetic-Instruct approach in enhancing the code generation capabilities of LLMs. The results for different sizes of the CodeLlama-Python model indicate that as the size of the base model increases, stronger performances can be achieved on most benchmarks. It should also be noted that HumanEval+ and MBPP+ are more trustworthy benchmarks as they are more robust to incorrect codes.
The results of GPT-4-Turbo and Mixtral 8x7B are also reported as references to illustrate the performance gap between OpenAI’s model and Mixtral. We used the Mixtral 8X7B as the generation model, which is a permissive model, whereas other approaches like WaveCoder or WizardCoder have used OpenAI’s models, which are not viable for the open-source community. To ensure a fair comparison between Genetic-Instruct and other synthetic data generation methods, we conducted more controlled experiments in Section 4.3.
It should be noted that any involvement of OpenAI’s models (Brown et al., 2020) in data synthesis has a marked positive effect on the downstream model, enabling substantial improvements in all models via data distillation. It can be seen that for WaveCoder and WizardCoder, results on all benchmarks are significantly improved compared to their base models. However, Genetic-Instruct surpasses these 7B models on EvalPlus datasets, indicating that such data distillation is not particularly robust. These results further suggest that while the choice of generator LLM is crucial in improving the results of downstream models, larger datasets can also provide more robust results.
An interesting observation of CodeLlama-70B’s results is its ability to surpass the Mixtral-8x7B model scores on all benchmarks despite using the relatively weaker Mixtral-8x7B models to generate data. The CodeLlama-70B-Python model surpasses Mixtral-8x7B’s scores on all benchmarks and shows further improvements after training on data generated by Mixtral-8x7B. This indicates that relatively weaker models can generate sufficiently diverse and challenging data, which, when applied to stronger domain-specific models, can outperform both. Essentially, we observe that improvements transfer from a relatively weaker model to a stronger model. This goes against the general trend of data distillation, which is usually from a stronger model to a weaker model. This result may enable an iterative self-improvement loop, such that the model fine-tuned on Genetic-Instruct data can be used to generate another dataset with higher quality for subsequent generations of models, similar to Noisy Student Training proposed by Xie et al. (2020).
@@ 4.3 Ablation Study
In this ablation study, we evaluate the effectiveness of the operations used in Genetic-Instruct on the performance of the generated samples. We perform three data generation ablations, labeled Crossover Only, Mutation Only, and Genetic-Instruct, which represent data generated using the Genetic-Instruct algorithm but with one task disabled (either crossover or mutation, respectively). We generate 250K samples for each dataset, and then uniformly sample without replacement subsets of 2K and 75K from this total dataset. We then train a CodeLlama-Python-7B model on each dataset and evaluate their performance. This ablation provides an objective comparison of Genetic-Instruct generation with alternatives Self-Instruct (Wang et al., 2023) and WizardCoder (Luo et al., 2024), which use crossover only and mutation only, respectively. By controlling variables such as compute, amount of data, generation model, and base model, we demonstrate a head-to-head comparison of generation techniques.
As seen in Table 2, crossover provides substantial improvements over the base model compared to mutation alone; however, both are necessary to achieve the best results. At just 2K samples, there is no clear winner between the three approaches, potentially due to the noisy sampling of 2K samples out of 250K, but all approaches improve on the baseline. As the dataset size increases to 75K and 250K, the combination of both crossover and mutation surpasses either separately, and this trend continues at larger data scales as measured on the EvalPlus benchmark.
Table 4: The effect of the population size on the performance of the Genetic-Instruct.
Dataset Size HumanEval HumanEval+ MBPP MBPP+
- 38.4 35.2 * 47.6 40.5 *
2K 46.3 41.5 56.4 45.4
75K 49.4 45.7 62.7 52.9
250K 51.8 47.6 63.9 54.6
405K 50.6 47.6 65.2 55.9
900K 51.2 48.2 64.9 56.4
4.4 Influence of the Generator Model on the Performance
A critical component of the Genetic-Instruct algorithm is the choice of the Instructor-LLM, Coder-LLM, and Judge-LLM. A powerful Instructor-LLM can generate more diverse instructions, a stronger Coder-LLM can produce code that is more likely to be correct and free from syntactic errors, and an effective Judge-LLM can better discern the quality of samples.
In Table 3, we utilized three different models (Mixtral 8x7B, Mixtral 8x22B, and Llama-3 70B) to generate a large amount of data in two to three generations and fine-tuned two base DeepSeek Coder models on them for comparison. We found that the quality of the final model strongly correlates with the quality of the data generated by it, and that different dataset sizes can achieve similar scores when two sufficiently powerful models are compared. An important observation is that while Mixtral 8x22B data shows strong results, obtaining similar scores to Llama 3 70B generated data in fewer samples, it required three generations of data to achieve similar scores to two generations for Llama 3 70B, thereby requiring significant additional compute.
Furthermore, when scaling up the dataset, the gap between the original and EvalPlus benchmarks is substantially reduced, as seen in the CodeLlama-7B experiments. Similarly, the results on CodeLlama-70B show a marginal improvement on the original HumanEval dataset but a substantial improvement on the EvalPlus counterpart. This indicates that despite saturation on the original benchmarks, training on more synthetic data improves the robustness of the models on extended benchmarks.
Table 5: Performance of the dataset generated by Genetic-Instruct with different seed datasets. CodeLlama-Python 7B has been used as the base model in the experiments.
Seed Dataset Size HumanEval HumanEval+ MBPP MBPP+
No SFT - 38.4 35.2 * 47.6 40.5 *
Code Alpaca 20K 49.4 45.7 63.4 53.4
Tiger-LeetCode 2K 50.6 47.6 65.2 55.9
4.5 Effectiveness of Scaling up the Population
In this section, we examine the impact of the size of generated samples on model performance. We generated 250K samples with Mixtral 8x7b and then uniformly sampled without replacement subsets of 2K and 75K from it. Subsequently, we trained a CodeLlama-Python 7B model on each dataset and evaluated their performance. The results are shown in Table 4. As the dataset size increases, performance generally improves, indicating that scaling synthetic data generation with our proposed algorithm can be beneficial. This trend is more evident in the MBPP+ and HumanEval+ datasets, showing that more synthetic data generated by Genetic-Instruct can enhance the quality and correctness of the codes the model generates.
4.6 Effect of the Initial Population
In this section, we evaluate the Genetic-Instruct algorithm on different datasets as the initial population to demonstrate the effectiveness of the proposed approach. We utilize CodeLlama-7B and two seed datasets: Tiger-leetcode and Code Alpaca333sahil2801/CodeAlpaca-20k. We generated roughly 400K samples with Genetic-Instruct using Code Alpaca as the seed set for comparison against our Tiger-leetcode baseline. Code Alpaca is an instruction dataset of around 20K samples created by ChatGPT 3.5. Table 5 shows the results of different seed datasets when used with Genetic-Instruct.
We observe that our method is robust to the seed dataset, capable of generating instruction data of similar quality for both seed sets, and significantly surpasses the base model’s results. There is a modest gap between the two seed datasets, which we attribute to the simplicity of instructions in Code Alpaca. It is important to note that Genetic-Instruct does not utilize the code solutions of the seed dataset when creating new instructions via the Instructor-LLM. We conjecture that the relative difficulty and complexity of Tiger-Leetcode’s questions lead to a more diverse synthetic dataset.
Given the similar performance observed between the Code Alpaca seeds generated by ChatGPT-3.5 and Tiger-leetcode, it appears that the success of the approaches utilizing OpenAI’s models is primarily due to the model used for generating instructions and codes, rather than the seeds themselves. This observation underscores the importance of the quality of LLMs employed in the generation process.
@@ 5 Conclusion
In this paper, we introduced Genetic-Instruct, a novel algorithm inspired by evolutionary principles, designed to generate large-scale synthetic coding instructions for large language models (LLMs). The algorithm utilizes two evolutionary operations, crossover and mutation, applied iteratively to a population pool to produce new instruction samples until the desired size is reached. Genetic-Instruct is specifically engineered to support parallel generation, making it a scalable solution for synthetic data creation. We benchmarked our approach against several baseline methods, demonstrating that Genetic-Instruct can generate instructions that lead to higher accuracy on code generation benchmarks. We also showed that increasing the size of the generated samples improves the performance of the final models. Furthermore, we evaluated the algorithm’s performance across various initial populations and base models, showcasing its robustness with respect to seed instructions and base LLMs.
@@ Appendix B Crossover Prompt
Cross-over Prompt You are asked to come up with a set of 20 diverse code generation task instructions. These task instructions will be given to a GPT model and we will evaluate the GPT model for completing the instructions. Here are the requirements: 1. Try not to repeat the verb for each instruction to maximize diversity. 2. The language used for the instruction also should be diverse. For example, you should combine questions with imperative instructions. 3. The type of instructions should be diverse. The list should include diverse types of programming tasks like open-ended generation, classification, editing, optimization etc. 4. A GPT language model should be able to complete the instruction. 5. The instructions should be in English. 6. The instructions should at least 1 to 2 sentences long. Either an imperative sentence or a question is permitted. 7. You should generate an appropriate input to the instruction. The input field should contain a specific example provided for the instruction. It should involve realistic data and should not contain simple placeholders. The input should provide substantial content to make the instruction challenging but should ideally not exceed 100 words. 8. Not all instructions require input. For example, when a instruction asks about some general information, "write a program to load a file.", it is not necessary to provide a specific context. In this case, we simply put "⟨noinput⟩" in the input field. 9. The output should be an appropriate response to the instruction and the input. 10. All tasks should be coding or programming-related. List of 20 tasks:
Few-Shot Examples ### 1. Instruction: Convert a Binary Search Tree to a sorted Circular Doubly-Linked List in place. You can think of the left and right pointers as synonymous to the predecessor and successor pointers in a doubly-linked list. For a circular doubly linked list, the predecessor of the first element is the last element, and the successor of the last element is the first element. We want to do the transformation in place. After the transformation, the left pointer of the tree node should point to its predecessor, and the right pointer should point to its successor. You should return the pointer to the smallest element of the linked list. 1. Input: root = 4,2,5,1,3 ### 2. Instruction: $\cdots$ $\cdots$ ### 4. Instruction:
Figure 3: Prompt template for the cross-over operation with few-Shot in-context learning
Appendix C Prompt Used for Generating Code with Coder-LLM

### v2

@@ 4.1 Experimental Settings
We used the AdamW optimizer (Kingma and Ba, 2015) for all supervised fine-tuning (SFT) experiments, with a learning rate of 5e-6 decaying to 5e-7 over three epochs, following a cosine annealing schedule Loshchilov and Hutter (2022). All models were trained using tensor parallelism and BF16 precision to accelerate the training process. Experiments were conducted using the NeMo framework Harper et al. (2025) and NeMo Aligner Shen et al. (2025).
@@
For high-throughput inference with large effective batch sizes, we used vLLM (Kwon et al., 2023) as the inference engine. Nucleus sampling (Holtzman et al., 2020) was employed for decoding, with a temperature of 1.2 for Instructor-LLM, and 1.0 for both Coder-LLM and Judge-LLM. To improve GPU utilization and speed up generation, we ran 20 colonies in parallel for each generation step. A maximum sequence length of 1024 tokens was set across all LLMs to optimize generation speed and memory usage.
@@
For Genetic-Instruct, the mutation probability ($M_{p}$) was set to 0.5 by default. During the mutation operation, a batch size of 100 ($B_{m}$) was used, while the crossover operation used a batch size of 10 ($B_{c}$). These values were chosen based on our observation that, the model generates approximately 10 unique instructions per generation on average, aiming to maintain a consistent number of generated samples per batch. In the crossover operation, Instructor-LLM used 3-shot in-context learning and was prompted to generate up to 20 new instructions.
Figure 2: The accuracy of Llama-3.1-8B trained on different data sizes. Code accuracy is calculated as the average of the model’s accuracy on all the four benchmarks. With scaling up the synthetic, accuracy improves but starts to show diminishing improvements later.
Gen. Algo. / Dataset $|S|$ MBPP MBPP+ HE HE+ Avg
Llama 3.1 8B Instruct - 73.0 62.7 66.5 61.6 65.9
Genetic Instruct 7.5M 79.9 69.1 66.5 63.4 69.7
Genetic Instruct 4M 76.5 66.9 65.9 62.8 68.0
Alternative Synthetic Data Generation Methods
WizardCoder 4M 72.8 62.4 65.9 61.6 65.7
Self-Instruct 4M 74.9 66.7 64.6 61.0 66.8
OSS-Instruct 4M 73.3 61.4 62.2 58.5 63.9
INVERSE-INSTRUCT 4M 59.8 49.2 29.3 26.2 41.1
Public Datasets
Code Parrot Apps 5k 39.7 34.7 29.9 28.1 33.1
TACO 25K 47.1 40.2 31.1 27.4 36.5
OpenCoder Stage 1 1M 67.2 57.1 66.5 61.0 62.9
OpenCoder Stage 2 170K 67.5 61.1 58.5 56.1 60.8
Code Alpaca 20K 31.8 26.7 24.4 20.7 25.9
Table 1: Comparison of Genetic-Instruct with other data generation algorithms and datasets. $|S|$ indicates numbers of training samples.
@@ 4.2 Performance Evaluation
In this section, we evaluate the effectiveness of our proposed approach for generating synthetic supervised fine-tuning (SFT) samples aimed at enhancing the coding capabilities of large language models (LLMs). We used Llama3.1-8B-Base (Grattafiori et al., 2024) as the base model and employed Mixtral-8x22B (Jiang et al., 2024) as the Instructor-LLM, Coder-LLM, and Judge-LLM.
Figure 2 illustrates the relationship between the size of the SFT dataset generated by Genetic-Instruct and coding accuracy. Coding accuracy is computed as the average model performance across all four benchmarks. We generated synthetic instructions across six generations, each consisting of approximately 1.5 million samples, totaling around 7.8 million samples. The results show a clear upward trend, where increasing the dataset size leads to significant improvements in accuracy. Notably, models trained on more than 3 million samples outperform the Llama3.1-8B-Instruct model. Starting from a baseline accuracy of approximately 45%, the Llama3.1-8B-Base model shows consistent improvement as the dataset grows, demonstrating the scalability and effectiveness of our synthetic data generation strategy. However, beyond approximately 6 million samples, the accuracy gains begin to plateau, indicating diminishing returns.
To show the effectiveness of Genetic-Instruct compared to other approaches, we evaluated the samples generated by Genetic-Instruct with some other baseline approaches which are designed for generating synthetic SFT data for coding problems. To make the comparisons fair, we re-implemented all the baseline approaches and performed the comparisons with the same generator model, seed population, base model for SFT, and size of training data. We did not rely on the results reported in the original papers, as each one used different generation models, seed populations, base models and benchmarks. Among these baselines, WizardCoder and Self-Instruct follow a similar paradigm to ours, using a collection of coding questions to expand into a larger instruction set. In contrast, OSS-Instruct Wei et al. (2024b) and INVERSE-INSTRUCT Wu et al. (2024) generate instructions from a large set of real code snippets.
For OSS-Instruct and INVERSE-INSTRUCT, we used around 1.4M Python functions extracted from Stack v2 Lozhkov et al. (2024) as the seed population, following the seed collection procedure adopted in Wei et al. (2024a), while for the rest of the baselines we used Tiger-Leetcode. The same number of samples are generated by each one of the approaches with three generations. Extra samples from the last generation are dropped randomly to make all the sizes exactly 4M. The results of 5 generations (7.5M) are also reported for Genetic-Instruct. We also evaluated some of the publicly available coding instruction datasets: Apps (Hendrycks et al., 2021), TACO (Li et al., 2023), and OpenCoder (Huang et al., 2024). All the results are presented in Table 1.
For OSS-Instruct and INVERSE-INSTRUCT, we used around 1.4M Python functions extracted from Stack v2 Lozhkov et al. (2024) as the seed population, following the procedure outlined in Wei et al. (2024a). For the remaining baselines, we used Tiger-Leetcode as the seed dataset. For each approach, we generated the same number of samples over three generations, and any extra samples from the final generation were randomly discarded to standardize the dataset size to 4 million. For Genetic-Instruct, we also report results with five generations (more than 7.5M samples). Additionally, we evaluated models fine-tuned on publicly available coding instruction datasets: Apps (Hendrycks et al., 2021), TACO (Li et al., 2023), and OpenCoder (Huang et al., 2024). The results are summarized in Table 1.
The results clearly highlight the superior performance of Genetic-Instruct across multiple evaluation metrics. Models trained on data generated by our method consistently outperform those trained on all baseline approaches and public datasets. In particular, our five-generation dataset achieves a significantly higher average accuracy of 69.7% compared to the best-performing public dataset, OpenCoder Stage 1, at 62.9%. Even our smaller dataset (4M) achieves an average of 68.0%, further underscoring the effectiveness and efficiency of our approach.
@@ 4.3 Ablation Study
In this ablation study, we assess the impact of mutation and crossover operations in Genetic-Instruct on the quality of generated data. We compare three setups: Crossover-Only, where only the crossover operation is used during data generation; Mutation-Only, where only the mutation operation is applied; and the full Genetic-Instruct approach, which employs both. For each, we generated three generations totaling 4 million samples and fine-tuned a Llama3.1-8B Base model to evaluate downstream performance. This setup allows us to assess the individual and combined impact of these genetic operators on downstream model performance. Mutation-Only resembles WizardCoder conceptually, but with a key distinction: it updates the evolving seed pool with newly generated samples, unlike WizardCoder, which evolves only the initial seeds.
As shown in Table 2, combining both operations yields the highest average accuracy across all benchmarks, confirming their complementary benefits. While Mutation-Only slightly outperforms the full approach on the HE benchmark, these findings suggest that while both operations individually contribute to improved performance, and their synergistic combination in Genetic-Instruct yields the most substantial overall gains in coding capability.
Generation Algorithm MBPP MBPP+ HE HE+ Avg
Cross-Over Only 74.9 66.7 64.6 61.0 66.8
Mutation Only 73.3 64.0 66.5 62.8 66.6
Genetic Instruct 76.5 66.9 65.9 62.8 68.0
Table 2: Comparing the effectiveness of different operations in the Genetic-Instruct algorithm. We generate 4 million samples for each experiment and used Llama 3.1 8B Base as the base model.
4.4 Influence of the Generator Model
Table 3 presents an ablation study evaluating the impact of different generator models on the quality of the synthetic data. We generated 1.5 million samples for each experiment with different generation models and then trained Llama3.1-8B-Base and Qwen2.5-7B-Base on them. The results indicate that the Qwen models Yang et al. (2024) outperform the Mixtral family across most benchmarks, highlighting that stronger LLMs tend to produce higher-quality synthetic data.
Base Model Generation Model MBPP MBPP+ HE HE+ Avg
Llama3.1 8B Mixtral 8x22B 72.8 64.0 62.8 59.8 64.8
Mixtral 8x7B 66.7 57.7 52.4 49.4 56.5
Qwen 32B 74.6 65.1 65.2 62.8 66.9
Qwen 7B 72.2 61.9 67.7 64.0 66.5
Qwen2.5 7B Mixtral 8x22B 79.1 67.2 72.6 65.9 71.2
Mixtral 8x7B 78.8 67.2 72.0 65.2 70.8
Qwen 32B 82.0 72.8 79.3 75.0 77.3
Qwen 7B 81.2 69.6 81.1 75.0 76.7
Table 3: Ablation study on the effect of the generator model on data generation quality.
Interestingly, Qwen-7B performs closely to Qwen-32B, suggesting that even a smaller model within the Qwen family is capable of generating high-quality training data. These findings imply that while the strength of the generator plays a key role in data quality, relatively smaller LLMs can still yield competitive performance, offering a more cost-effective alternative for synthetic data generation.
@@ 5 Conclusion
We introduced Genetic-Instruct, a novel algorithm inspired by evolutionary principles for generating synthetic coding instructions for large language models. Genetic-Instruct is specifically designed to support parallel generation, making it a highly scalable solution for synthetic data creation. We benchmarked our approach against several baseline methods and publicly available datasets, and the results consistently demonstrated its effectiveness in improving performance on downstream code generation tasks.
@@ Appendix B Crossover Prompt
Crossover Prompt You are asked to come up with a set of 20 diverse code generation task instructions. These task instructions will be given to a GPT model and we will evaluate the GPT model for completing the instructions. Here are the requirements: 1. Try not to repeat the verb for each instruction to maximize diversity. 2. The language used for the instruction also should be diverse. For example, you should combine questions with imperative instructions. 3. The type of instructions should be diverse. The list should include diverse types of programming tasks like open-ended generation, classification, editing, optimization etc. 4. A GPT language model should be able to complete the instruction. 5. The instructions should be in English. 6. The instructions should at least 1 to 2 sentences long. Either an imperative sentence or a question is permitted. 7. You should generate an appropriate input to the instruction. The input field should contain a specific example provided for the instruction. It should involve realistic data and should not contain simple placeholders. The input should provide substantial content to make the instruction challenging but should ideally not exceed 100 words. 8. Not all instructions require input. For example, when a instruction asks about some general information, "write a program to load a file.", it is not necessary to provide a specific context. In this case, we simply put "⟨noinput⟩" in the input field. 9. The output should be an appropriate response to the instruction and the input. 10. All tasks should be coding or programming-related. List of 20 tasks:
Few-Shot Examples ### 1. Instruction: Convert a Binary Search Tree to a sorted Circular Doubly-Linked List in place. You can think of the left and right pointers as synonymous to the predecessor and successor pointers in a doubly-linked list. For a circular doubly linked list, the predecessor of the first element is the last element, and the successor of the last element is the first element. We want to do the transformation in place. After the transformation, the left pointer of the tree node should point to its predecessor, and the right pointer should point to its successor. You should return the pointer to the smallest element of the linked list. 1. Input: root = 4,2,5,1,3 ### 2. Instruction: $\cdots$ $\cdots$ ### 3. Instruction:
Figure 4: Prompt template for the crossover operation with few-shot in-context learning
Appendix C Prompts for Coder-LLM